The half-life of a certain radioatcive material is 85 days. An initial amount of the material has a mass of 801 kg. Write an exponential function that models the decay of this material remains after 10 days.
Answer by ankor@dixie-net.com(15623)

The half-life of a certain radioatcive material is 85 days.
An initial amount of the material has a mass of 801 kg.
Write an exponential function that models the decay of this material remains after 10 days.
:
The half life formula: A = Ao*2^(-t/h)
where
Ao = initial amt (801)
A = resulting amt after t time elapses
t = time (10)
h = half-life of substance (85)
:
A = 801 * 2^(-10/85)